Dear sir!

Please solve this question.
Q 1. Write down an equation or equations for each of the following.
a. A plane parallel to xy-plane at a distance of 4 units from origin on negative z-axis.
b. A curve produced due to the intersection of a circular cylinder and yz-plane. Center of this circular cylinder is x-axis (that is, cylinder is around x-axis) and its radius is one.
c. A curve produced due to the intersection of a sphere, center at origin and radius 3, and xy-plane.
: Dear sir!

Please solve this question.
Q 1. Write down an equation or equations for each of the following.
a. A plane parallel to xy-plane at a distance of 4 units from origin on negative z-axis.
b. A curve produced due to the intersection of a circular cylinder and yz-plane. Center of this circular cylinder is x-axis (that is, cylinder is around x-axis) and its radius is one.
c. A curve produced due to the intersection of a sphere, center at origin and radius 3, and xy-plane.

Answer by Alan3354(1941)

Please solve this question.
Q 1. Write down an equation or equations for each of the following.
a. A plane parallel to xy-plane at a distance of 4 units from origin on negative z-axis.
--------------------
z = -4. For all values of x and y, z is 4.
--------------------
b. A curve produced due to the intersection of a circular cylinder and yz-plane. Center of this circular cylinder is x-axis (that is, cylinder is around x-axis) and its radius is one.
--------------------------------
The intersection is a circle about the origin.
y^2 + z^2 = 1
c. A curve produced due to the intersection of a sphere, center at origin and radius 3, and xy-plane.
----------------
This is also a circle about the origin. Any plane intersecting a sphere will be a circle.
x^2 + y^2 = 3^2
x^2 + y^2 = 9